# Reflection on Code Contributor Demographics and Collaboration Patterns in the Rust Community

### v2

@@ 4.2. RQ2: What are the collaboration patterns and interaction dynamics among Rust contributors?
The structural properties of these networks were further examined through the calculation of various graph metrics, as presented in Table 2. The reciprocity of the directed versions of these networks provides insight into the mutual nature of collaborations. Reciprocity measures the likelihood of node pairs having reciprocal connections—in this context, whether contributors review and submit PRs to each other. The Cargo project exhibited a reciprocity score of 0.5, indicating a relatively balanced two-way collaboration among contributors. In contrast, the Rust and Rust Analyzer projects showed lower reciprocity scores. All the undirected networks, constructed only using ISO Codes and member status nodes separately, displayed low density values, which is characteristic of real-world social networks and OSS networks in particular (Mezouar et al., 2019).
Figure 3. Rust Analyzer Project contributors social network
Here is the updated table with the values rounded to four decimal places:
Here’s the updated table with the values rounded to two decimal places:
@@
Figure 4. Cargo Project contributors social network
4.3. RQ3: To what extent do sub-communities and top contributors within the Rust ecosystem exhibit diversity?
We analyzed the sub-communities formed within the social undirected graphs of the Rust, Rust Analyzer, and Cargo projects. The diversity within these sub-communities was quantified using Blau’s Index, as presented in Table LABEL:tab:blau-index. Eight sub-communities were identified in the Rust project, with Blau’s Index scores ranging from 0.000 to 0.883. Most of these communities exhibited high geographical diversity, with Blau’s Index scores above 0.85, indicating a diverse mix of contributors from different countries. Community 1 and Community 2 had Blau’s Index scores of 0.857 and 0.883, respectively, reflecting substantial diversity. An exception was Community 8, which had a Blau’s Index of 0.000 due to comprising only two contributors from the same country, making it an outlier.
Two sub-communities were formed for the Rust Analyzer project, displaying high geographical diversity with Blau’s Index scores of 0.889 and 0.895. This result suggests that contributors within these communities are from a wide range of countries, contributing to a diverse collaborative environment.
In the Cargo project, six sub-communities were identified. Community 1 had a high Blau’s Index of 0.914, indicating significant geographical diversity. The remaining five communities each had a Blau’s Index of 0.500, suggesting moderate diversity within these groups.
To understand the influence and roles of influential contributors across sub-communities, we calculated centrality measures—degree centrality, betweenness centrality, and eigenvector centrality—for the contributors across the projects, as shown in Table 4. Across all projects, the top contributors identified through centrality measures were predominantly men, with the Likely Gender being men or Unknown. There were no women contributors among the most central nodes.
@@
The users with the most edges connected to them indicating top contributions via degree centrality metrics are shown in Table 4. Geographically, in the Rust project, they were primarily from the United States (US), Germany (DE), and France (FR). In the Rust Analyzer project, top contributors hailed from Germany (DE), Romania (RO), Japan (JP), and Iran (IR), indicating a broader geographical spread compared to the Rust project. The Cargo project featured top contributors from Taiwan (TW), the United States (US), and Germany (DE). The same geographic distribution was seen in Table 4 representing users connected to the most influential nodes in the respective project networks.
Table 4 depicts the top users with the highest contributions going through them(node) in the network, representing the best nodes for the fastest flow of contributions. Both members and non-members of the Rust GitHub organization were among the top contributors. For instance, in the Rust project, one of the top contributors with a high degree of centrality was a non-member. This suggests that significant contributions and influence within the projects are not confined to official members. The betweenness centrality scores varied across the projects. In the Rust project, the top betweenness centrality scores ranged from 0.10 to 0.26, indicating a less centralized network where contributions do not predominantly pass through a few individuals. Conversely, the Rust Analyzer project exhibited higher betweenness centrality scores, with the top contributor scoring 0.72, suggesting that specific contributors play critical roles as this indicates many contributions (PR authorship or reviews) go through them from/to different parts of the network acting as a central piece. Similarly, the top betweenness centrality scores in the Cargo project were 0.77 and 0.48, indicating a more centralized network structure.
@@
Ranking ISO Code Likely Gender Score Is Member ISO Code Likely Gender Score Is Member ISO Code Likely Gender Score Is Member
@@ 5. Discussion
Geographic disparities: While the survey indicates any form of user engagement or participation from around 100 countries, our findings from RQ1 indicate that active code contributions are concentrated in fewer countries across the studied projects. The Rust project shows contributions from 50 countries, Rust Analyzer from 36, and Cargo from 30, with 55 unique countries across projects. This suggests that active contributions are more concentrated while community engagement is widespread. The United States and Germany are prominent in both the survey and contributor data, signifying strong representation. However, some countries like China, which ranks third in the survey (6% of participants), have minimal representation among top contributors, appearing only in the Cargo project with a modest contribution share (0.93%). This finding aligns with previous research on OSS participation imbalances (Gasparini et al., 2020b), such as the 90-9-1 rule, which highlights that 90% of users are passive observers, 9% contribute occasionally, and only 1%—the “superusers”—generate the majority of contributions (Nielsen, 2006). Conversely, countries such as Romania and Iran, which are not highlighted in the survey’s top ten, have substantial contribution shares in the Rust Analyzer project (13.96% and 11.24%, respectively)—indicating that certain regions may have a small number of highly active contributors despite lower overall community participation.
@@ 5.2. Collaboration Patterns and Interaction Dynamics
Varying Contributor Dynamics: The central clustering of member nodes, particularly in the Rust and Rust Analyzer projects, indicates that members tend to collaborate closely with one another and may serve as hubs within the network. Their central positions suggest they are integral to the project’s development, possibly taking on code review, mentoring, and coordination roles. The dispersion of member nodes in the Cargo project may reflect different collaboration practices or a more decentralized organizational structure. Consequentially, the high reciprocity between all the nodes in the Cargo project compared to others is to be noticed. A reciprocity score of 0.5 in the Cargo project implies a higher degree of mutual collaboration, where contributors frequently engage in bidirectional interactions. This indicates equal collaboration dynamics, where contributors both author and review PRs reciprocally. In contrast, the lower reciprocity scores in the Rust and Rust Analyzer projects suggest a more unidirectional collaboration pattern. This may be due to hierarchical structures, where specific contributors predominantly take on reviewing roles, possibly due to their expertise or organizational positions, as observed in the network figures of the respective repositories with dense member hubs. This gives us an idea of how decentralized structures lead to more collaboration, measured through reciprocity in this context.
@@
5.3. Diversity Within Sub-Communities
Substantial Geographic Representation: The high Blau’s Index scores in most sub-communities indicate substantial geographical diversity, particularly in the Rust and Rust Analyzer projects. This diversity suggests that contributors from different countries are collaborating effectively within these projects.
@@ 5.3. Diversity Within Sub-Communities
Moderate Sustainability: The presence of members and non-members among the top contributors demonstrates the open and collaborative nature of the Rust ecosystem. Non-members can achieve significant influence and centrality within the network, indicating that official membership is not a prerequisite for impactful contributions. This openness can foster a more diverse and dynamic contributor base, as it lowers barriers to participation. The differences in betweenness centrality scores across the projects highlight variations in network structure and collaboration dynamics. The Rust project’s lower betweenness centrality suggests a more distributed network, spreading influence across multiple contributors and reducing dependency on specific individuals. In contrast, the higher betweenness centrality in the Rust Analyzer and Cargo projects indicates that certain contributors act as key connectors or gatekeepers. This centralization may pose risks if these individuals reduce their participation, potentially impacting the project’s progress and collaboration flow. The geographic diversity among top contributors, especially in the Rust Analyzer project, reflects the project’s ability to attract and integrate contributors from a variety of countries, including those less represented in the Rust project. Contributors from Romania, Japan, and Iran play significant roles, suggesting that the project has effective mechanisms for engaging and supporting international contributors. Overall, while the Rust ecosystem shows substantial geographical diversity, particularly within sub-communities, there is a clear need to improve gender diversity among contributors. Addressing gender disparities can enhance inclusivity, impact, and innovation.
@@ 6.1. Limitations
While our study provides valuable insights into the collaboration patterns and diversity within the Rust community, some limitations must be acknowledged. Firstly, the constraints of the GitHub API posed significant challenges. The API restricts the amount and type of data that can be programmatically extracted, limiting access to detailed user profiles and historical contribution records beyond the specified time frame. As a result, our dataset may lack comprehensive information that could offer a more nuanced understanding of contributor behaviors and demographics. This limitation hinders our ability to capture the full spectrum of factors influencing collaboration within the community.
While NamSor is a reputable tool used in previous studies (IEEE, 2023; Zhang et al., 2024), it may not always accurately reflect individuals’ self-identified genders, especially in cases involving gender-neutral names or culturally diverse naming conventions. We recognize that a considerable portion of the contributors had unspecified geographical locations and gender data, which imposed constraints on our findings and analysis. Nevertheless, we adhered to publicly accessible information about the contributors to preserve their privacy and ensure the research adhered to ethical standards. Additionally, we successfully obtained data for the majority of the top contributors across all three evaluated repositories.
The study excludes a small number of self-loops from the network graphs that have been created to focus on collaboration. The metrics we used, such as reciprocity, Blau’s Index, and centrality measures, are based solely on data from a one-year period. While this provides a snapshot of recent collaboration patterns, it does not capture long-term trends or account for contributors whose activity levels fluctuate over time. Thus, our findings reflect only a specific timeframe and may not be indicative of historical dynamics within the Rust community.
Our findings are limited to the Rust ecosystem and may not be generalizable to other OSS communities or projects in different domains. The unique characteristics of Rust’s development practices, community culture, and organizational structure could influence the collaboration patterns observed, necessitating caution when extrapolating these results to other contexts.
@@ 6.2. Future Work
There are several avenues that could be pursued to deepen our understanding of collaboration and diversity in OSS communities. Conducting longitudinal studies over multiple years (Schueller et al., 2022) would provide insights into how collaboration patterns and contributor demographics evolve over time, helping to identify trends in contributor retention, shifts in community engagement, and the impact of organizational changes. Building on the insights from this quantitative analysis, foundations supporting OSS communities, such as the Rust Foundation,   [^12] https://foundation.rust-lang.org/ should implement targeted initiatives to enhance sustainability and inclusivity within the community. These initiatives may involve lowering entry barriers for new developers, establishing mentorship programs, and providing resources to support diverse demographic groups. By leveraging collaboration patterns and diversity metrics, the Rust Foundation can create a more welcoming environment that encourages participation from underrepresented regions and genders, ensuring the community’s growth and resilience. Furthermore, distributing responsibilities more evenly among contributors can mitigate network centralization and reduce the risk of burnout, promoting a healthier and more balanced contributor base. It is needed particularly when leading tech firms are cutting back on their DEI initiatives.  [^13] https://www.bbc.com/news/articles/cgmy7xpw3pyo
Expanding the scope and methodology undertaken in this research to include multiple OSS communities across different domains and levels of organizational maturity would allow for comparative analyses. Such studies could identify common patterns or unique differences, contributing to a broader understanding of collaboration dynamics in OSS development.
Finally, integrating additional network metrics or analytical frameworks could offer a more nuanced view of the collaboration networks. For example, analyzing motifs, clustering coefficients, or applying machine learning algorithms might uncover deeper insights into the structural properties of the networks.
@@ 7. Conclusion
This study investigated the demographic diversity and collaboration dynamics within the Rust programming language community by analyzing PR data from three pivotal repositories: Rust (the core compiler), Rust Analyzer (an IDE front-end), and Cargo (the package manager). Our analysis revealed that active contributors in these projects exhibit lower geographic and gender diversity compared to the broader Rust user community, as indicated by the 2023 Rust Annual Survey.
Specifically, most top contributors are men, and contributions are predominantly concentrated in a few countries. This concentration highlights a disparity between the global user base and the active contributor community, suggesting potential barriers to broader participation. The social network analysis highlighted the importance of non-members contributing substantially despite not being official Rust GitHub members. However, the reliance on a small number of central individuals, particularly in Rust Analyzer and Cargo, poses potential risks to project resilience. The diversity within sub-communities, demonstrated substantial geographic diversity in most clusters, particularly within the Rust and Rust Analyzer projects. Nonetheless, gender diversity remains significantly low, with very few women contributors among the most central and influential nodes.
@@ 8. Data availability
We have made our anonymized data and scripts publicly available (Anonymous, 2024) to ensure reproducibility and facilitate replication of our findings.

### v3

@@ 4.2. RQ2: What are the collaboration patterns and interaction dynamics among Rust contributors?
The structural properties of these networks were further examined through the calculation of various graph metrics, as presented in Table 2. The reciprocity of the directed versions of these networks provides insight into the mutual nature of collaborations. Reciprocity measures the likelihood of node pairs having reciprocal connections, in this context, whether contributors review and submit PRs to each other. The Cargo project exhibited a reciprocity score of 0.5, indicating a relatively balanced two-way collaboration among contributors. In contrast, the Rust and Rust Analyzer projects showed lower reciprocity scores. All the undirected networks, constructed only using ISO Codes and member status nodes separately, displayed low density values, which is characteristic of real-world social networks and OSS networks in particular (Mezouar et al., 2019).
Figure 3. Rust Analyzer Project contributors’ social network
@@
Figure 4. Cargo Project contributors’ social network
4.3. RQ3: To what extent do subcommunities and top contributors within the Rust ecosystem exhibit diversity?
We analyzed the subcommunities formed within the social undirected graphs of the Rust, Rust Analyzer, and Cargo projects. The diversity within these subcommunities was quantified using Blau’s Index, as presented in Table LABEL:tab:blau-index. Eight subcommunities were identified in the Rust project, with Blau’s Index scores ranging from 0.000 to 0.883. Most of these communities exhibited high geographical diversity, with Blau’s Index scores above 0.85, indicating a diverse mix of contributors from different countries. Community 1 and Community 2 had Blau’s Index scores of 0.857 and 0.883, respectively, reflecting substantial diversity. An exception was Community 8, which had a Blau’s Index of 0.000 due to comprising only two contributors from the same country, making it an outlier.
Two subcommunities were formed for the Rust Analyzer project, displaying high geographical diversity with Blau’s Index scores of 0.889 and 0.895. This result suggests that contributors within these communities are from a wide range of countries, contributing to a diverse collaborative environment.
In the Cargo project, six subcommunities were identified. Community 1 had a high Blau’s Index of 0.914, indicating significant geographical diversity. The remaining five communities each had a Blau’s Index of 0.500, suggesting moderate diversity within these groups.
@@
To understand the influence and roles of influential contributors across subcommunities, we calculated centrality measures, degree centrality, betweenness centrality, and eigenvector centrality, for the contributors across the projects, as shown in Table 4. Across all projects, the top contributors identified through centrality measures were predominantly men, with the Likely Gender being men or Unknown. There were no women contributors among the most central nodes. The users with the most edges connected to them, indicating top contributions via degree centrality metrics, are shown in Table 4. Geographically, in the Rust project, they were primarily from the United States (US), Germany (DE), and France (FR). In the Rust Analyzer project, top contributors hailed from Germany (DE), Romania (RO), Japan (JP), and Iran (IR), indicating a broader geographical spread compared to the Rust project. The Cargo project featured top contributors from Taiwan (TW), the United States (US), and Germany (DE). The same geographic distribution was seen in the third part of 4, representing users connected to the most influential nodes in the respective project networks, captured through eigenvector centrality.
The middle part of Table 4 depicts the top users with the highest contributions going through them (node) in the network, representing the best nodes for the fastest flow of contributions. Both members and non-members of the Rust GitHub organization were among the top contributors. For instance, in the Rust project, one of the top contributors with a high degree of centrality was a non-member. This suggests that significant contributions and influence within the projects are not confined to official members. The betweenness centrality scores varied across the projects. In the Rust project, the top betweenness centrality scores ranged from 0.10 to 0.26, indicating a less centralized network where contributions do not predominantly pass through a few individuals. Conversely, the Rust Analyzer project exhibited higher betweenness centrality scores, with the top contributor scoring 0.72, suggesting that specific contributors play critical roles as this indicates many contributions (PR authorship or reviews) go through them from/to different parts of the network acting as a central piece. Similarly, the top betweenness centrality scores in the Cargo project were 0.77 and 0.48, indicating a more centralized network structure.
@@
Rank ISO Gender Score Member ISO Gender Score Member ISO Gender Score Member
@@
  • Few entries in bold depict contributors whose country or membership status are outliers.
@@ 5. Discussion
Geographic Disparities: While the survey indicates any form of user engagement or participation from around 100 countries, our findings from RQ1 indicate that active code contributions are concentrated in fewer countries across the studied projects. The Rust project shows contributions from 50 countries, Rust Analyzer from 36, and Cargo from 30, with 55 unique countries across projects. This suggests that active contributions are more concentrated while community engagement is widespread. The United States and Germany are prominent in both the survey and contributor data, signifying strong representation. However, some countries like China, which ranks third in the survey (6% of participants), have minimal representation among top contributors, appearing only in the Cargo project with a modest contribution share (0.93%). This finding aligns with previous research on OSS participation imbalances (Gasparini et al., 2020b), such as the 90-9-1 rule, which highlights that 90% of users are passive observers, 9% contribute occasionally, and only 1%, the “superusers”, generate the majority of contributions (Nielsen, 2006). Conversely, countries such as Romania and Iran, which are not highlighted in the survey’s top ten, have substantial contribution shares in the Rust Analyzer project (13.96% and 11.24%, respectively), indicating that certain regions may have a small number of highly active contributors despite lower overall community participation.
@@ 5.2. Collaboration Patterns and Interaction Dynamics
Varying Contributor Dynamics: The central clustering of member nodes, particularly in the Rust and Rust Analyzer projects, indicates that members tend to collaborate closely with one another and may serve as hubs within the network. Their central positions suggest they are integral to the project’s development, possibly taking on code review, mentoring, and coordination roles. The dispersion of member nodes in the Cargo project may reflect different collaboration practices or a more decentralized organizational structure. Consequently, the high reciprocity between all the nodes in the Cargo project compared to others is to be noticed. A reciprocity score of 0.5 in the Cargo project implies a higher degree of mutual collaboration, where contributors frequently engage in bidirectional interactions. This indicates equal collaboration dynamics, where contributors both author and review PRs reciprocally. In contrast, the lower reciprocity scores in the Rust and Rust Analyzer projects suggest a more unidirectional collaboration pattern. This may be due to hierarchical structures, where specific contributors predominantly take on reviewing roles, possibly due to their expertise or organizational positions, as observed in the network figures of the respective repositories with dense member hubs. This gives us an idea of how decentralized structures lead to more collaboration, measured through reciprocity in this context.
@@
5.3. Diversity Within Subcommunities
Substantial Geographic Representation: The high Blau’s Index scores in most subcommunities indicate substantial geographical diversity, particularly in the Rust and Rust Analyzer projects. This diversity suggests that contributors from different countries are collaborating effectively within these projects.
@@ 5.3. Diversity Within Subcommunities
Moderate Sustainability: The presence of members and non-members among the top contributors demonstrates the open and collaborative nature of the Rust ecosystem. Non-members can achieve significant influence and centrality within the network, indicating that official membership is not a prerequisite for impactful contributions. This openness can foster a more diverse and dynamic contributor base, as it lowers barriers to participation. The differences in betweenness centrality scores across the projects highlight variations in network structure and collaboration dynamics. The Rust project’s lower betweenness centrality suggests a more distributed network, spreading influence across multiple contributors and reducing dependency on specific individuals. In contrast, the higher betweenness centrality in the Rust Analyzer and Cargo projects indicates that certain contributors act as key connectors or gatekeepers/blockers. This centralization, although showing a group of vital contributors pivotal for the ecosystem, may pose risks if these individuals reduce their participation, potentially impacting the project’s progress and collaboration flow. The geographic diversity among top contributors, especially in the Rust Analyzer project, reflects the project’s ability to attract and integrate contributors from a variety of countries, including those less represented in the Rust project. Contributors from Romania, Japan, and Iran play significant roles, suggesting that the project has effective mechanisms for engaging and supporting international contributors. Overall, while the Rust ecosystem shows substantial geographical diversity, particularly within subcommunities, there is a clear need to improve gender diversity among contributors. Addressing gender disparities is likely to enhance inclusivity, impact, and innovation.
@@ 6.1. Limitations
While our study provides valuable insights into collaboration and diversity in the Rust community, several limitations must be acknowledged. First, the GitHub API imposes limits on the granularity of extractable data, limiting access to detailed user profiles and complete historical records beyond the chosen one-year window. Additionally, strict API rate limits significantly delayed data collection, requiring multiple iterations to capture all relevant pull request information.
Location information was based on self-reported GitHub profile entries and likely reflects contributors’ current locations rather than their nationalities, which may limit the accuracy of capturing broader ethnic or racial diversity. Automated tools such as NamSor, while commonly used in prior studies (IEEE, 2023; Zhang et al., 2024), may also misclassify gender, particularly for gender-neutral or culturally diverse names. Similarly, the geolocation parsing through NLTK introduced reliability concerns, and manual corrections during the verification process carried a risk of human error.
Our mining and analysis scripts, developed in Python and R, could be subject to undetected coding errors, despite careful validation. Furthermore, we excluded self-loops in the constructed collaboration graphs to better focus on contributor interactions, which may slightly influence network density metrics. All network and diversity metrics were derived from a single year of data, offering a contemporary but time-limited view of collaboration patterns.
Finally, the findings are specific to the Rust ecosystem and should be generalized to other OSS communities with caution. The exclusive reliance on quantitative pull request data, without complementary qualitative insights such as interviews or surveys, may leave out important contextual factors influencing collaboration dynamics.
@@ 6.2. Future Work
Several avenues could be pursued to address these limitations and deepen our understanding of collaboration and diversity in OSS communities. Conducting longitudinal studies across multiple years (Schueller et al., 2022) would help reveal trends in contributor retention, shifts in engagement, and the effects of organizational or ecosystem changes over time. Building on the insights from this study, organizations like the Rust Foundation [^12] https://foundation.rust-lang.org/ could implement targeted initiatives to improve sustainability and inclusivity, such as lowering entry barriers, creating mentorship programs, and providing support for underrepresented groups. These efforts are especially important as many tech firms scale back diversity initiatives. [^13] https://www.bbc.com/news/articles/cgmy7xpw3pyo Expanding this research to include multiple OSS communities across different domains and maturity levels would allow comparative analyses, highlighting common collaboration patterns or unique ecosystem-specific challenges. Additionally, integrating further network metrics and analytical frameworks, such as analyzing motifs, clustering coefficients, or applying machine learning techniques, could offer a deeper structural understanding of collaboration networks. Finally, complementing quantitative analyses with qualitative research, through surveys or interviews with contributors and maintainers, could provide critical context behind observed patterns, offering a more comprehensive view of the social and organizational factors shaping OSS collaboration.
@@ 7. Conclusion
This study investigated the demographic diversity and collaboration dynamics within the Rust programming language community by analyzing PR data from three pivotal repositories: Rust (the core compiler), Rust Analyzer (an IDE front-end), and Cargo (the package manager). Our analysis revealed that active contributors to these projects exhibit lower geographic and gender diversity compared to the broader Rust user community, as indicated by the 2023 Rust Annual Survey.
Specifically, most of the top contributors are men, and contributions are predominantly concentrated in a few countries. This concentration highlights a disparity between the global user base and the active contributor community, suggesting potential barriers to broader participation. The social network analysis highlighted the importance of non-members contributing substantially despite not being official Rust GitHub members. However, the reliance on a small number of central individuals, particularly in Rust Analyzer and Cargo, poses potential risks to project resilience. The diversity within subcommunities demonstrated substantial geographic diversity in most clusters, particularly within the Rust and Rust Analyzer projects. Nonetheless, gender diversity remains significantly low, with very few women contributors among the most central and influential nodes.
@@ 8. Data availability
We have made our anonymized data and scripts publicly available (Dandamudi and Rodríguez-Pérez, 2024) to ensure reproducibility and facilitate replication of our findings.
Acknowledgements.
This research was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), Grant RGPIN-2022-03265.